# Knowledge Editing in Language Models via Adapted Direct Preference Optimization

### v1

@@ 3.5 Optimization Step
Comparison with FT: KDPO enables more reliable KE, by allowing the term-cancellation phenomenon mentioned above. Additionally, because KDPO makes relative adjustments based on a reference model $\pi_{ref}$, it prevents the model from deviating significantly from the original weights. Furthermore, KDPO allows for controllable editing capabilities through the use of the $\beta$ parameter. Empirical evidence of this can be seen in Tab. 1, where the average locality of KDPO is significantly higher than that of FT-M and FT-L (which are fine-tuning variations) for the three models examined.
@@ 5 Results
This section presents our results across multiple dataset, LLMs, and KE methods. We start by showing the results of our multiple edits study. Then, we demonstrate our method’s ability to handle general, non KE, LLM benchmarks. Lastly, we show an extensive ablation study. Additionally, prompt examples and empirical edit results are presented in Section 4. The prompts were selected from the ZsRE dataset to give some intuition on prompt types and edit challenges. We also show comparative responses from multiple KE methods shown in this paper. Some baselines did not converge in all cases and thus were not added to the results. Those issues and other baseline implementation details are discussed in Section  B.1 of the supplementary material.Further results appear in the supplementary material Section D. Our method uses the same hyperparameters for all models (size ranging from 0.5B-8B) and datasets, while we optimized the hyperparameters of baseline methods to achieve competitive results.
@@ 5.2 Knowledge Editing in Small Language Model
Different sizes of language models vary in the way they train, predict, and react to KE. We compare our method to other method as well as to standard DPO. Our method shows very promising results for smaller models like Qwen1.5-0.5B. Fig. 4 shows the average results on three leading datasets . Results are presented in the supplementary material Section D, and shows once again that KDPO keeps opens a gap against baseline method on 500 sequential edits.
@@ 7 Limitations
The primary limitation of our method, which is inherited from DPO, is the need to keep a copy of the model, leading to an increase in the memory footprint. However, it is important to note that recent works, such as those by [Meng et al., 2024] and [Azar et al., 2024], are actively addressing this challenge, offering potential solutions to reduce this hurdle.
@@ Appendix B Implementation Details
We used the same hyperparameters for all models across all datasets. We optimized using the Adam optimizer with a learning rate of $41e-$ 4, number of cycles, $n$, equal to 10, and the number of steps in each cycle is 8. We only optimized layer 21 as done in FT-M. However, empirically examined, the average number of cycles is 4 and the average number of total steps is 29. All experiments were done using PyTorch [Paszke et al., 2019] on Nvidia A100 GPU.
@@ B.1 Baseline implementation
For the baselines, we used the hyperparameters in the EasyEdit repo; However not all methods contains hyperparameter for all the examined models; In those cases we make grid search on several hyperparameters. For example, FT-M edits layer 31, but Qwen1.5-0.5B has only 24 layers, thus we ran with editing layers 13-21, layer 15 was the best and their result presented. Further, FT-M mostly use normalization factor of 5e-5, using this factor result in a very low edit success, after searching, we’ve used a factor of 5e-4. Some of the methods we were not able to reach decent results for example PMET on Llama2 didn’t converge to appropriate numbers, and on the other large model we constantly had Memory issues. For MEMMIT we were able to reach results Llama but not on other model due to memory issues.
Appendix C Example Prompts Edit Results
Tab. 4 shows example prompts from the ZsRE dataset.
Appendix D Further Results
D.1 Small Language Model Results
The results for KE on Qwen1.5-0.5B are present in Tab. 5. We also show a radar plot of the 500 sequential edits in Fig. 4
@@ D.1 Small Language Model Results
D.2 Ablation Study Detailed results

### v2

@@ 3.5 Optimization Step
Comparison with FT: KDPO enables more reliable KE, by allowing the term-cancellation phenomenon mentioned above. Additionally, because KDPO makes relative adjustments based on a reference model $\pi_{ref}$, it leads to model edits that are more local. Furthermore, KDPO allows for controllable editing capabilities through the use of the $\beta$ parameter. Empirical evidence of this can be seen in Tab. 1, where the average locality of KDPO is significantly higher than that of FT-M and FT-L (which are fine-tuning variations) for the three models examined.
@@ 5 Results
This section presents our results across multiple dataset, LLMs, and KE methods. We start by showing the results of our multiple edits study. Then, we demonstrate our method’s ability to handle general, non KE, LLM benchmarks. Lastly, we show an extensive ablation study. Additionally, prompt examples and empirical edit results are presented in Tab. 2. The prompts were selected from the ZsRE dataset to give some intuition on prompt types and edit challenges. We also show comparative responses from multiple KE methods shown in this paper. Some baselines did not converge in all cases and thus were not added to the results. Those issues and other baseline implementation details are discussed in Section  B.1 of the supplementary material. Further results appear in the supplementary material Section C. Our method uses the same hyperparameters for all models (size ranging from 0.5B-8B) and datasets, while we optimized the hyperparameters of baseline methods to achieve competitive results.
@@ 5.2 Knowledge Editing in Small Language Model
Different sizes of language models vary in the way they train, predict, and react to KE. We compare our method to other method as well as to standard DPO. Our method shows very promising results for smaller models like Qwen1.5-0.5B. Fig. 4 shows the average results on three leading datasets . Results are presented in the supplementary material Section C, and shows once again that KDPO keeps opens a gap against baseline method on 500 sequential edits.
@@ 7 Limitations
The primary limitation of our method, which is inherited from DPO, is the need to keep a copy of the model, leading to an increase in the memory footprint. However, it is important to note that recent works, such as those by Meng et al. [2024] and Azar et al. [2024], are actively addressing this challenge, offering potential solutions to reduce this hurdle.
Acknowledgments
This work was supported by a grant from the Tel Aviv University Center for AI and Data Science (TAD).
@@ Appendix B Implementation Details
We used the same hyperparameters for all models across all datasets. We optimized using the Adam optimizer with a learning rate of $41e-$ 4, number of cycles, $n=$ 10, and the number of steps in each cycle is $s=$ 8. We only optimized layer 21 as done in FT-M. However, empirically examined, the average number of cycles is 4 and the average number of total steps is 29. All experiments were done using PyTorch [Paszke et al., 2019] on Nvidia A100 GPU.
@@ B.1 Baseline implementation
For the baselines, we used the hyperparameters in the EasyEdit repository. However not all methods contain hyperparameter for all the examined models; In those cases we conduct a grid search on several hyperparameters. For example, FT-M usually edits layer 31, but Qwen1.5-0.5B has only 24 layers, thus we sweep layers 13-21, layer 15 was the best and that result was presented in the paper. Further, FT-M mostly use normalization factor of 5e-5, using this factor results in a very poor "edit success" in some models, after searching for better factors, we’ve used a factor of 5e-4 (only for the models that got poor results using 5e-5). For some methods we were not able to reach decent results. For example, PMET on LLaMA2-7B did not converge to appropriate numbers, and on the other LLMs we constantly had GPU memory issues. For MEMIT, we were able to reach results on LLaMA2-7B but not on other model due to GPU memory issues. SERAC and MEND we were not able to get reasonable results in training on Qwen1.5 models and LLaMA3-8B.
Appendix C Further Results
C.1 Small Language Model Results
The results for KE on Qwen1.5-0.5B are present in Tab. 5. We also show a radar plot of the 500 sequential edits in Fig. 4. The results shows that the effectiveness gap of our method increases as the number of edits increases, this also can be deduced by other experiments in this paper.
@@ C.1 Small Language Model Results
C.2 Ablation Study Detailed results
@@
Method Training Batch Editing Weight Update IKE Yes No None AdaLoRA Yes Yes LoRA ROME Yes No Partial FT-L Yes Yes Full FT-M Yes Yes Partial MEMIT Yes Yes Partial DPO Yes Yes Partial Ours Yes Yes Partial
Table 6: Comparison of Different Method’s Properties